Determine whether the following relation is reflexive, symmetric, antisymmetric and/or transitive.

(x, y) ∈ R if x ≥ y, where R is the set of positive integers.


1
Expert's answer
2021-12-01T17:40:50-0500

Let us determine whether the relation "(x, y) \u2208 R"  if "x \u2265 y" is reflexive, symmetric, antisymmetric and/or transitive.

Since "x\ge x" for each positive integer "x," we conclude that the relation is reflexive.

Taking into account that "(2,1)\in R" but "(1,2)\notin R," we conclude that the relation is not symmetric.

If "(x,y)\in R" and "(y,x)\in R," then "x\ge y" and "y\ge x." It follows that "x\ge y\ge x," and hence  "y=x."We conclude that the relation is antisymmetric.

Since "(x,y)\in R" and "(y,z)\in R" imply "x\ge y" and "y\ge z," we conclude that "x\ge y\ge z", and thus  "x\ge z." It follows that "(x,z)\in R," and thus the relation is transitive.